A test is composed of six multiple choice questions where each question has 4 choices. If the answer choices for each question are equally likely, find the probability of answering 3 OR 4 questions correctly
Answer by stanbon(57387)

A test is composed of six multiple choice questions where each question has 4 choices. If the answer choices for each question are equally likely, find the probability of answering 3 OR 4 questions correctly
-----------
Binomial Problem with n=6 and p = 1/4
----
P(3 or 4) = P(3) + P(4)
==
6C3*(1/4)^3*(3/4)3 + 6C4*(1/4)^4*(3/4)^2
=========================================
= 0.1318 + 0.0330
======================
Cheers,
Stan H.